find three consecutive odd integers such that three times the third is seventeen more than the sum of the first and second one.

Answer by stanbon(57337)

find three consecutive odd integers such that three times the third is seventeen more than the sum of the first and second one.
-----
1st: 2x-1
2nd: 2x+1
3rd: 2x+3
----
Equation:
3(2x+3) = 4x+17
6x+9 = 4x+17
2x = 8
x = 4
-----
1st: 2x-1 = 7
2nd: 9
3rd: 11
=================
Cheers,
Stan H.